2. B (25.0 m, 30° counterclockwise from the vertical). Find the x- and y-componen the coordinate systems shown. of B in each of 30°
Answer #1

here,

a)

the x-component of B , Bx = 25 * cos(25) m

Bx = 22.7 m

the y-component of B , By = 25 * sin(25) m

By = 10.57 m

b)

when axis are rotated theta = 30 degree counterclockwise

the x-component of B , Bx = 25 * cos(25 - theta) m

Bx = 25 * cos(25 - 30) m

Bx = 24.9 m

the y-component of B , By = 25 * sin(25 - theta) m

By = 25 * sin(25 - 30) m

By = - 2.18 m